OpenStudy (anonymous):

Solve 25x2 – 64 = 0 {64/25} {8/5} {-64/25, 64/25} {-8/5, 8/5}

OpenStudy (btaylor):

add 64 to both sides, then divide by 25. you get x^2=64/25. Take the positive and negative square roots, and you get {-8/5, 8/5}

OpenStudy (maheshmeghwal9):

use\[a^2-b^2=(a+b)(a-b).\]